The Harmonic Oscillator frequency For Oxygen nucleus is approximately how much times of Harmonic Oscillatory frequency for Barium nucleus?
1
Expert's answer
2020-06-11T10:35:45-0400

It is hard to tell without numerical values, but you can use the following expression to estimate is:


"\frac{\omega_\text{O}}{\omega_\text{B}}=\sqrt{\frac{k_\text{O}m_{R\text{B}}}{k_{\text{B}}m_{R\text{O}}}},"

where "k" and "m_R" are bond force constants and reduced masses for oxygen or barium.